# 2D-TPE: Two-Dimensional Positional Encoding Enhances Table Understanding for Large Language Models

### v2

@@ 1. Introduction
Tables are highly structured and rich in information, making them indispensable and widely used in the real world. From financial reports to scientific data, tables serve as an efficient means of organizing and presenting complex relationships and patterns. As large language models (LLMs) continue to advance (Achiam et al., 2023; Anil et al., 2023; Touvron et al., 2023), and the interest in developing LLM-based agents for completing specific tasks grows (Tang et al., [n. d.]; Zhang et al., 2024d; Liu et al., [n. d.]; Guan et al., 2024; Cheng et al., 2024), endowing LLMs with the ability to accurately comprehend and reason over tabular data has emerged as a crucial research direction (Ye et al., 2023; Zhang et al., 2024b, c).
@@
A fundamental challenge for LLM-based table understanding lies in the inherent mismatch between the two-dimensional (2D) structure of tables and the one-dimensional (1D) input format required by LLMs. To bridge the gap, existing methods typically flatten the tabular data into a sequence of tokens (Zhang et al., 2024c). While this simple approach offers a straightforward way to adapt tabular data to existing LLMs, it disregards the spatial relationships and contextual information encoded within the layout of tables. Consequently, LLMs may struggle to perform accurate analysis and reasoning over tabular data, even for seemingly simple tasks. For example, we devise two proxy tasks (as illustrated in Figure 1), namely Counting-Stars and Locating-Values, to assess the capability of LLMs to identify specific cells based on their positional relations to another cell, which is a crucial foundation for table understanding (more details of the tasks are presented in §3.3). Empirical evaluations demonstrate that LLMs equipped with conventional 1D positional encodings perform remarkably poorly on these tasks, achieving an accuracy of less than 5% and 20% in $20\times 20$ tables, respectively. The results highlight the detrimental impact of flattening operations on preserving table structures and underscore the need for more effective encoding methods to facilitate LLMs’ perception of tabular data.
@@
In this work, we explore a simple yet effective approach to representing tabular data in LLMs, by which both semantic information and spatial information can be well captured. The core idea is inspired by recent advances in the development of large multimodal models. In general terms, tables and images are similar in some sense, as both modalities convey information through points distributed across a two-dimensional area, making spatial information critically important (Lu et al., 2024). Hence, encouraged by the remarkable success of positional encoding techniques in vision-language models (VLMs), we propose a novel approach termed “2D-TPE” (Two-Dimensional Table Positional Encoding) to address the challenge of tabular structure representation in LLMs. Unlike traditional VLMs that encode patch positions using 2D coordinates—a method less suitable for tables where cells may contain multiple tokens, 2D-TPE employs a more flexible approach. In a nutshell, 2D-TPE enables each attention head to dynamically select a permutation order for perceiving the context, where each permutation represents a distinct traversal mode for the table, such as column-wise or row-wise traversal. Through dynamic permutation selection, our approach allows for flexible and adaptive context perception, enabling the model to explore various traversal patterns and capture the most relevant spatial dependencies. This adaptability is particularly valuable in scenarios where the importance of specific dimensions or relationships within the table may vary, ensuring that the model can effectively focus on the most salient aspects of the data. In this way, 2D-TPE can mitigate the risk of losing essential spatial information while maintaining computational efficiency.
@@
We conduct experiments with an open-source LLM on five benchmarks, covering a wide range of table understanding tasks, including question-answering, type annotation, relation extraction, and entity linking. The evaluation results consistently demonstrate the superiority of 2D-TPE over strong baselines employing the same LLM but with conventional 1D positional encoding, with most improvements (3 out of 5 tasks) exhibiting statistical significance (Sign-test, $p$-value $<0.05$). Notably, 2D-TPE exhibits exceptional robustness and scalability when confronted with tables of varying sizes, maintaining stable performance even when the table quadruples. These findings underscore the substantial potential of 2D-TPE in tackling real-world challenges involving large-scale tabular data. In stark contrast, the performance of 1D positional encoding deteriorates significantly as table sizes increase, highlighting their fragility in handling complex table structures. Remarkably, 2D-TPE achieves an excellent balance between efficacy and efficiency. Compared to vanilla Transformers, the additional computational cost in terms of TFLOPs and memory usage is negligible, with an increase of less than 2%. Furthermore, the inference time per example only experiences a modest 13% increase. In summary, 2D-TPE paves the way for more effective and versatile table understanding systems.
@@
II. We introduce a powerful and versatile positional encoding method “2D-TPE,” which enables LLMs to dynamically select different permutation orders for perceiving the table’s context, effectively preserving the spatial relationships within the tabular data.
@@ 2.1. LLM-based Table Understanding
To tailor LLMs for table-related tasks, several studies have curated specialized tabular datasets for instruction-tuning purposes (Jiang et al., 2022; Liu et al., 2022). For instance, Table-GPT (Li et al., 2023) synthesized diverse instruction-completion pairs from real tables. TableLlama (Zhang et al., 2024b) introduced the comprehensive TableInstruct dataset, encompassing a diverse array of table-related tasks, alongside a base model that exhibits remarkable generalization capabilities across multifaceted benchmarks.
@@
LLMs have shown a remarkable reasoning capacity (Wei et al., 2022a) through the Chain of Thought (CoT) prompting strategy (Wei et al., 2022b; Chen, 2023) that solves complex queries step by step. Consequently, considerable research efforts have been dedicated to developing various prompting techniques to improve LLM performance in table understanding tasks. For example, Dater (Ye et al., 2023) prompted the LLM to extract a sub-table containing the most pertinent information related to the question and decompose the question into a series of sub-questions. TaCo (Zheng et al., 2023) incorporates the CoT approach into small-scale tabular LMs, specifically to address mathematical questions involving tabular data. PROTRIX (Wu and Feng, 2024) presents a Plan-then-Reason framework that strategically plans the problem-solving pathway and engages in step-by-step reasoning.
@@
To address the structured nature and inherent logic of tabular data, several studies have explored the integration of LLMs with auxiliary tools such as SQL and Python interpreters, enabling precise calculations and location-based operations. The text-to-SQL paradigm (Nan et al., 2023) provides a direct method for translating natural language queries into executable SQL statements, facilitating the retrieval and manipulation of data stored in tabular formats. Furthermore, Binder (Cheng et al., [n. d.]) introduced the incorporation of Python tools to handle complex computations and detailed cell positioning within tables. ReAcTable (Zhang et al., 2024a) and Chain-of-Table (Wang et al., 2024) proposed interleaving the reasoning and tool-invocation steps, allowing LLMs to dynamically invoke appropriate tools at each step of the problem-solving process.
@@ 2.2. Table Modeling
Recent studies attempted to adapt attention mechanisms to accommodate the inherently 2D structure of tabular data. TABERT (Yin et al., 2020) implemented column-wise self-attention on the output of row-wise self-attention, thereby enhancing each token’s awareness of its positional information. TURL (Deng et al., 2022) and MATE (Eisenschlos et al., 2021) proposed a constrained attention mechanism, wherein each token attends exclusively to tokens within the same row or column. TABLEFORMER (Yang et al., 2022) incorporated learnable bias terms into attention scores between tokens, based on their relative positional relationships in the table. In contrast to these approaches that rely on dedicated models to capture table structures, 2D-TPE offers a more effective solution and aligns better with mainstream LLMs by encoding the spatial relationships within the standard self-attention mechanism.
@@ 2.3.1. Absolute Positional Encoding (APE)
One intuitive approach is to map position indices into learnable embeddings, as employed in the BERT (Devlin et al., 2019) and GPT (Radford et al., 2018). However, this method fails to generalize to positions that have not been encountered during training, leading to substantial performance degradation when the inference length exceeds the training length (Su et al., 2024). To address the challenge, Vaswani et al. (2017) introduced the sinusoidal position embeddings that maps a position index $m$ to a fixed embedding $\boldsymbol{P}_{m}$ through a series of sinusoidal functions. Under this formulation, for any position offset $\Delta_{m}$, the positional embedding $\boldsymbol{P}_{m+\Delta_{m}}$ for the token $m+\Delta m$ can be represented as a linear function of $\boldsymbol{P}_{m}$, thereby facilitating the model’s potentials to generalize patterns based on relative positions.
@@
Prior research also attempted to encode 2D tabular data using learnable embeddings. Among the efforts, TAPAS (Herzig et al., 2020) assigned each table token multiple positional embeddings as inputs, explicitly denoting its row and column indices. TABBIE (Iida et al., 2021) integrated the outputs of two distinct Transformer models, each equipped with independent positional embeddings, to separately encode the row and column context of each token. To deal with hierarchical tables, TUTA (Wang et al., 2021) introduced tree-based position embeddings for hierarchical positional information of each token. Unlike 2D-TPE, the above methods with learnable positional embeddings still potentially face the challenges of length extrapolation.
@@ 2.3.2. Relative Positional Encoding (RPE)
RPE focuses on inter-token relative distances, enhancing the model’s length extrapolation capability. The most commonly employed RPE techniques are ALiBi (Press et al., 2022) and RoPE(Su et al., 2024), both of which are applied on every self-attention layer without additional trainable parameters. ALiBi introduced a linear bias to each attention term. RoPE modulates the query and key vectors using rotary matrices derived from absolute position indices, with the attention weights remaining solely contingent on the relative positional offset between the query and key.
@@
2D RPE has been recently explored in image encoding since images also possess an inherent 2D structure, necessitating the use of 2D positional encodings when decomposed into patch sequences. Unified-IO-2 (Lu et al., 2024) extends RoPE to 2D by splitting the query and key vectors of attention heads and applying separate rotary embeddings constructed from horizontal and vertical coordinates to the respective halves. While this approach has demonstrated efficacy in improving performance on various image understanding tasks (Lu et al., [n. d.]; Heo et al., 2024), it is not directly applicable to tables due to two key challenges: (1) It cannot distinguish multiple distinct tokens within the same cell sharing identical 2D coordinates; and (2) It processes attention for the two dimensions independently, potentially hindering the model’s ability to learn inter-token positional patterns in tables, as different tokens can be assigned the same horizontal or vertical coordinates. This circumstance is less severe for image encoding since positional patterns are relatively fixed when splitting images into a fixed number of patches. In contrast, 2D-TPE does not rely on 2D coordinates but instead leverages different permutation orders of tokens to capture structural information in tables. In this way, 2D-TPE effectively addresses the aforementioned challenges and can be scaled to incorporate more permutation orders to capture more structural information beyond horizontal and vertical directions. While the implementation of 2D-TPE in this paper is based on RoPE, it is readily adaptable to other RPE techniques (e.g., ALiBi).
@@ 4.2. Training Objective
Furthermore, in order to encourage the model to select a specific permutation order for each attention head and each token more distinctly, we introduce an auxiliary loss to sharpen the distribution of router weights by minimizing its entropy (Chen et al., 2023):
@@ 4.3. Candidate Permutation Orders
Using proper permutation orders as candidates in 2D-TPE is a crucial consideration. Let us first investigate tokens within the table. One can traverse a table following different orders, each inducing distinct position indices and representing varying inductive biases regarding the proximity of cells within the table. Common traversal modes include row-wise, column-wise, diagonal, Hilbert-curve (Hilbert and Hilbert, 1935), and Z-order-curve (Dugundji, 1989) traversals. In this paper, we illustrate the effect of 2D-TPE using two representative traversal modes to obtain the permutation orders (i.e., $J=2$ in Equation 7): row-wise and column-wise traversals, both proceeding from top-left to bottom-right. This choice can be readily extended to accommodate other traversal modes. Note that the relative distances between tokens in the same cell (e.g., “United” and “_States” in Figure 2) always remain the same regardless of permutation orders.
@@ 5.1.1. Evaluation Benchmarks
To rigorously evaluate the performance of 2D-TPE, we conduct experiments on five diverse table understanding tasks, as summarized in Table 2. We curate all datasets from TableInstruct (Zhang et al., 2024b) and maintain only those examples with correct table structures and lengths not exceeding 4,096 tokens counted using the MiniCPM tokenizer (Hu et al., 2024).
@@ 5.1.2. Baselines
We compare 2D-TPE against several aforementioned strong baselines that employ the conventional 1D RoPE for positional encodings, including Row-wise Traversal, Column-wise Traversal, and Constrained Attention.
@@ 5.1.3. Implementation Details
We implement 2D-TPE and baselines by fine-tuning MiniCPM-2B-SFT (Hu et al., 2024) on each dataset, setting the hyper-parameter $\lambda$ in Eq. 15 to 1, the batch size to 64, the learning rate to 2e-5, the length limits to 4,096 and the warm-up steps to 3% of 2 epochs. Appendix B further describes the influence of hyper-parameter settings in detail. We employ DeepSpeed with ZeRO-2 (Rajbhandari et al., 2020) and Flash-attention-2 (Dao, 2024) for all methods, except Constrained Attention that modifies the attention mask, and thus becomes incompatible with Flash-Attention. We use greedy decoding for inference.

### v3

@@ 1. Introduction
Tables are highly structured and rich in information, making them indispensable and widely used in the real world. From financial reports to scientific data, tables serve as an efficient means of organizing and presenting complex relationships and patterns. As large language models (LLMs) continue to advance (Achiam et al., 2023; Anil et al., 2023; Touvron et al., 2023), and the interest in developing LLM-based agents for completing specific tasks grows (Tang et al., [n. d.]; Zhang et al., 2024e; Liu et al., [n. d.]; Guan et al., 2024; Cheng et al., 2024), endowing LLMs with the ability to accurately comprehend and reason over tabular data has emerged as a crucial research direction (Ye et al., 2023; Zhang et al., 2024c, d).
@@
A fundamental challenge for LLM-based table understanding lies in the inherent mismatch between the two-dimensional (2D) structure of tables and the one-dimensional (1D) input format required by LLMs. To bridge the gap, existing methods typically flatten the tabular data into a sequence of tokens (Zhang et al., 2024d). While this simple approach offers a straightforward way to adapt tabular data to existing LLMs, it disregards the spatial relationships and contextual information encoded within the layout of tables. Consequently, LLMs may struggle to perform accurate analysis and reasoning over tabular data, even for seemingly simple tasks. For example, we devise two proxy tasks (as illustrated in Figure 1), namely Counting-Stars and Locating-Values, to assess the capability of LLMs to identify specific cells based on their positional relations to another cell, which is a crucial foundation for table understanding (more details of the tasks are presented in §3.3). Empirical evaluations demonstrate that LLMs equipped with conventional 1D positional encodings perform remarkably poorly on these tasks, achieving an accuracy of less than 5% and 20% in $20\times 20$ tables, respectively. The results highlight the detrimental impact of flattening operations on preserving table structures and underscore the need for more effective encoding methods to facilitate LLMs’ perception of tabular data.
@@
In this work, we present a novel positional encoding approach, dubbed “2D-TPE” (Two-Dimensional Table Positional Encoding), designed to effectively capture both semantic and spatial information inherent in 2D tabular data while seamlessly accommodating 1D textual data. Akin to tables, images also convey information through points distributed across a 2D space, rendering spatial information critically important (Lu et al., 2024). However, recent advancements in vision-language models (VLMs) employing 2D positional encoding (Wang et al., 2024a) are not readily transferable to the table understanding domain. This is because tables exhibit a dynamic nature with varying sizes and variable-length tokens within individual cells, in contrast to the fixed-size patches used in image representations. Consequently, we argue that tables should be treated as a unique modality, distinct from both textual and image data, to fully leverage their inherent structure and spatial relationships. In a nutshell, 2D-TPE enables each attention head to dynamically select a permutation order for perceiving the context, where each permutation represents a distinct traversal mode for the table, such as column-wise or row-wise traversal. Through dynamic permutation selection, our approach allows for flexible and adaptive context perception, enabling the model to explore various traversal patterns and capture the most relevant spatial dependencies. This adaptability is particularly valuable in scenarios where the importance of specific dimensions or relationships within the table may vary, ensuring that the model can effectively focus on the most salient aspects of the data. In this way, 2D-TPE can mitigate the risk of losing essential spatial information while maintaining computational efficiency.
@@
We conduct experiments with an open-source LLM on five benchmarks, covering a wide range of table understanding tasks, including question-answering, type annotation, relation extraction, and entity linking. The evaluation results consistently demonstrate the superiority of 2D-TPE over strong baselines employing the same LLM, with most improvements (3 out of 5 tasks) exhibiting statistical significance (Sign-test, $p$-value $<0.05$). Notably, 2D-TPE exhibits exceptional robustness and scalability when confronted with tables of varying sizes, maintaining stable performance even when the table quadruples. These findings underscore the substantial potential of 2D-TPE in tackling real-world challenges involving large-scale tabular data. In stark contrast, the performance of 1D positional encoding deteriorates significantly as table sizes increase, highlighting their fragility in handling complex table structures. Remarkably, 2D-TPE achieves an excellent balance between efficacy and efficiency. Compared to vanilla Transformers, the additional computational cost in terms of TFLOPs and memory usage is negligible, with an increase of less than 2%. Furthermore, the inference time per example only experiences a modest 13% increase. In summary, 2D-TPE paves the way for more effective and versatile table understanding systems.
@@
II. We introduce a versatile positional encoding method “2D-TPE,” which enables LLMs to dynamically select different permutation orders for perceiving the table’s context, efficiently and effectively preserving the spatial relationships within the tabular data.
@@ 2.1. LLM-based Table Understanding
To tailor LLMs for table-related tasks, several studies have curated specialized tabular datasets for instruction-tuning purposes (Jiang et al., 2022; Liu et al., 2022). For instance, Table-GPT (Li et al., 2023) synthesized diverse instruction-completion pairs from real tables. TableLlama (Zhang et al., 2024c) introduced the comprehensive TableInstruct dataset, supporting varied tasks and showcasing a model with broad generalization across benchmarks.
@@
LLMs have shown a remarkable reasoning capacity (Wei et al., 2022a) through the Chain of Thought (CoT) prompting strategy (Wei et al., 2022b; Chen, 2023) that solves complex queries step by step. Consequently, considerable research efforts have been dedicated to developing various prompting techniques to improve LLM performance in table understanding tasks. For example, Dater (Ye et al., 2023) prompted the LLM to extract key sub-tables and decompose questions into sub-questions. TaCo (Zheng et al., 2023) used the CoT approach in tabular LMs for mathematical queries. PROTRIX (Wu and Feng, 2024) introduced a Plan-then-Reason framework for structured problem-solving and step-by-step reasoning.
@@
To address the structured nature and inherent logic of tabular data, several studies have explored the integration of LLMs with auxiliary tools such as SQL and Python interpreters, enabling precise calculations and location-based operations. The text-to-SQL paradigm (Nan et al., 2023) translated natural language queries into SQL for data retrieval and manipulation. Binder (Cheng et al., [n. d.]) integrated Python tools for complex computations and precise cell positioning in tables. ReAcTable (Zhang et al., 2024b) and Chain-of-Table (Wang et al., 2024b) interleaved reasoning and tool invocation, enabling LLMs to dynamically use tools throughout the problem-solving process.
@@ 2.2. Table Modeling
HyTrel (Chen et al., 2023a) transformed tabular data into hypergraphs to capture structural attributes but lacked compatibility with mainstream LLMs. Recent studies have adapted attention mechanisms to accommodate the inherently 2D structure of tabular data. TABERT (Yin et al., 2020) layered column-wise self-attention on top of row-wise self-attention, enhancing positional awareness among tokens. StruBERT (Trabelsi et al., 2022) employed a combination of horizontal and vertical self-attention. TURL (Deng et al., 2022) and MATE (Eisenschlos et al., 2021) restricted attention to tokens within the same row or column. TABLEFORMER (Yang et al., 2022) introduced learnable biases to adjust attention scores based on token positions. Unlike these models, 2D-TPE efficiently encodes spatial relationships within the standard self-attention framework, better aligning with mainstream LLMs.
@@ 2.3.1. Absolute Positional Encoding (APE)
One intuitive approach is to map position indices into learnable embeddings, as employed in the BERT (Devlin et al., 2019) and GPT (Radford et al., 2018). However, this method fails to generalize to positions that have not been encountered during training, leading to substantial performance degradation when the inference length exceeds the training length (Su et al., 2024). To address the challenge, Vaswani et al. (2017) introduced the sinusoidal position embeddings that mapped a position index $m$ to a fixed embedding $\boldsymbol{P}_{m}$ through a series of sinusoidal functions. Under this formulation, for any position offset $\Delta_{m}$, the positional embedding $\boldsymbol{P}_{m+\Delta_{m}}$ for the token $m+\Delta m$ can be represented as a linear function of $\boldsymbol{P}_{m}$, thereby facilitating the model’s potentials to generalize patterns based on relative positions.
@@
Prior research also attempted to encode 2D tabular data using learnable embeddings. Among the efforts, TAPAS (Herzig et al., 2020) used multiple positional embeddings per table token to denote row and column indices. TABBIE (Iida et al., 2021) combined outputs from two Transformers with unique positional embeddings to encode row and column contexts. TUTA (Wang et al., 2021) introduced tree-based embeddings for hierarchical table positions. Unlike 2D-TPE, the above methods with learnable positional embeddings still potentially face the challenges of length extrapolation.
@@ 2.3.2. Relative Positional Encoding (RPE)
RPE focuses on inter-token relative distances, enhancing the model’s length extrapolation capability. The most commonly employed RPE techniques are ALiBi (Press et al., 2022) and RoPE(Su et al., 2024), both of which are applied on every self-attention layer without additional trainable parameters. ALiBi introduced a linear bias to each attention term. RoPE modulated the query and key vectors using rotary matrices derived from absolute position indices, with the attention weights remaining solely contingent on the relative positional offset between the query and key.
@@
2D RPE has been adapted for image encoding due to images’ inherent 2D structure, requiring positional encodings for patch sequences. Unified-IO-2 (Lu et al., 2024) adapted RoPE to 2D by dividing the query and key vectors of attention heads and applying separate rotary embeddings from horizontal and vertical coordinates. Although effective for image tasks (Lu et al., [n. d.]; Heo et al., 2024), its application to tables is limited by inability to distinguish tokens within the same cell and independent dimensional attention processing, which may miss inter-token positional patterns. This issue is less critical in fixed-patch image encoding. In contrast, 2D-TPE uses varied permutation orders of tokens to capture structural table information, overcoming these issues and scaling to incorporate more permutation orders to capture more structural information beyond horizontal and vertical directions. While 2D-TPE in this study builds on RoPE, it can adapt to other RPE techniques (e.g., ALiBi).
@@ 4.2. Training Objective
Furthermore, in order to encourage the model to select a specific permutation order for each attention head and each token more distinctly, we introduce an auxiliary loss to sharpen the distribution of router weights by minimizing its entropy (Chen et al., 2023b):
@@ 4.3. Candidate Permutation Orders
Using proper permutation orders as candidates in 2D-TPE is a crucial consideration. Let us first investigate tokens within the table. One can traverse a table following different orders, such as row-wise, column-wise, diagonal, Hilbert-curve (Hilbert and Hilbert, 1935), and Z-order-curve (Dugundji, 1989) traversals, each of which induces distinct position indices and representing varying inductive biases regarding the proximity of cells within the table. In this paper, we illustrate the effect of 2D-TPE using two representative traversal modes to obtain the permutation orders (i.e., $J=2$ in Equation 7): row-wise and column-wise traversals, both proceeding from top-left to bottom-right. This choice can be readily extended to accommodate other traversal modes, which is left for future work. Note that the relative distances between tokens in the same cell (e.g., “United” and “_States” in Figure 2) always remain the same regardless of permutation orders.
@@ 5.1.1. Evaluation Benchmarks
To rigorously evaluate the performance of 2D-TPE, we conduct experiments on five diverse table understanding tasks, as summarized in Table 2. We curate all datasets from TableInstruct (Zhang et al., 2024c) and maintain only those examples with correct table structures and lengths not exceeding 4,096 tokens counted using the MiniCPM tokenizer (Hu et al., 2024).
@@ 5.1.2. Baselines
We compare 2D-TPE against several aforementioned strong baselines using conventional 1D RoPE, including Row-wise Traversal, Column-wise Traversal, and Constrained Attention. Additionally, we compare 2D-TPE with TABBIE (Iida et al., 2021), which integrates row and column embeddings for enhanced table representations, and Multimodal Rotary Position Embedding (M-RoPE), originally used in Qwen2-VL (Wang et al., 2024a) to jointly capture image and text positional information. We adapt M-RoPE to tabular data by decomposing the RoPE positional embeddings into three independent components for row, column, and cell positions.
@@ 5.1.3. Implementation Details
We train TABBIE on each dataset according to the settings described in (Iida et al., 2021) and implement 2D-TPE and other baselines by fine-tuning MiniCPM-2B-SFT (Hu et al., 2024) on each dataset, selected for its impressive performance (Zhang et al., 2024a) and ease of industrial deployment. We set the hyper-parameter $\lambda$ in Eq. 15 to 1, the batch size to 64, the learning rate to 2e-5, the length limits to 4,096 and the warm-up steps to 3% of 2 epochs. Appendix B further describes the influence of hyper-parameter settings in detail. We employ DeepSpeed with ZeRO-2 (Rajbhandari et al., 2020) and Flash-attention-2 (Dao, 2024) for all methods, except Constrained Attention that modifies the attention mask, and thus becomes incompatible with Flash-Attention. We use greedy decoding for inference.